OpenStudy (anonymous):

Explain how you could use either multiplication or division to solve the same equation.

OpenStudy (anonymous):

Suppose you have the equation: 10x = 50 Divide both sides by 10, \[\frac{10x}{10} = \frac{50}{10}\] x = 5

OpenStudy (radar):

It is the same operation, multiplication by 2, is same as dividing by 1/2. Provide an equation

OpenStudy (anonymous):

And for multiplication: Suppose you have the equation like: \[\frac{x}{3} = 9\] Multiply with 3 both the sides: \[3 \times \frac{x}{3} = 9 \times 3\] x = 27

OpenStudy (anonymous):

thanks

OpenStudy (radar):

Do you have a specific equation in mind?

OpenStudy (anonymous):

no it was a question given to me on an assignment

OpenStudy (radar):

O.K.